### v1

# SeedProteo: Accurate De Novo All-Atom Design of Protein Binders

Abstract
We present SeedProteo, a diffusion-based model for de novo all-atom protein design. We demonstrate how to repurpose a cutting-edge folding architecture into a powerful generative design framework by effectively integrating self-conditioning features. Extensive benchmarks highlight the model’s capabilities across two distinct tasks: in unconditional generation, SeedProteo exhibits superior length generalization and structural diversity, maintaining robustness for long sequences and complex topologies; in binder design, it achieves state-of-the-art performance among open-source methods, attaining the highest in-silico design success rates, structural diversity and novelty.
[…]
Figure 1: Comparison of binder design success and diversity across ten target proteins. We benchmark exclusively against open-source methods using SeedFold in-silico metrics.
1 Introduction
Proteins perform a vast array of functions within organisms. As the evolution process spans millions to billions of years, the field of protein design aims to automate and accelerate it through computational methods. Since the function of proteins is confered via the structure, which is determined by the sequence, a line of research focuses on generating sequences [dauparasRobustDeepLearning2022, zhengStructureinformedLanguageModels2023, alamdariProteinGenerationEvolutionary2024, wang2024dplm] and structures [yimSE3DiffusionModel2023b, watsonNovoDesignProtein2023, ingrahamIlluminatingProteinSpace2023, zhang2025odesign]. Recently, a new paradigm has emerged emphasizing the co-generation of sequence and structure [wang2025dplm2, hsieh2025dplm2_1, chen2025apm, multiflow]. However, improving sample diversity while preserving the coherence between modalities, specifically sequence-structure consistency, remains a formidable challenge. To address this, recent works have pivoted toward all-atom modeling [quPallatomUnlockingNew2025, geffnerLaProteinaAtomisticProtein2025]. By directly designing proteins at the all-atom level, these methods naturally derive amino acid identities from atomic coordinates, achieving superior consistency. These methodological advancements have significantly propelled the field of de novo protein design, particularly in binder design, the creation of mini-binders for specific therapeutic targets [zambaldi2024novo]. This domain has witnessed a rapid evolution, progressing from classical hallucination-based methods [pacesaOneshotDesignFunctional2025] or backbone-based design model [zhang2025odesign, ren2025pxdesign] to the current wave of all-atom diffusion models [butcher2025novo, chai2025zero, team2025latent], all striving for enhanced designability and binding affinity.
In this study, we introduce SeedProteo, a diffusion-based model that learns the manifold of proteins at the all-atom level. SeedProteo mimics the architecture of AlphaFold3 [abramsonAccurateStructurePrediction2024, wohlwendBoltz1DemocratizingBiomolecular2024, passaroAccurateEfficientBinding], which has proven its effectiveness in protein folding. Through unifying the representation of different amino acids, we significantly enhance its generative power when the protein sequence and homology are not provided. A major issue in all-atom modeling is the backbone-sidechain inconsistency (i.e., the sequence-structure inconsistency), which means that the sidechain atoms are locally plausible but the derived sequence fails to fold into the global backbone structure. SeedProteo iteratively evolves the sequence and structure and integrates two components to enhance the modeling ability of two modalities: (i) SeedProteo enhances the structural feasibility through the integration of geometric reasoning in the denoising process. To be specific, we reuse the denoised structure from the previous timestep [chen2022analog] to stabilize the sampling process and conduct geometric reasoning in the pair representation space. (ii) SeedProteo conducts energy minimization within the sequence space modeled by a Markov random field (MRF) [metzler2005markov]. By integrating high-order couplings among distinct positions, we are able to sample globally feasible sequences and prevent being trapped in the local minima of simple patterns constituted by high-frequency amino acids. We conduct experiments on two tasks: unconditional sampling and binder design. The results demonstrate that:
1) In unconditional generation, SeedProteo exhibits superior length generalization and structural diversity compared to state-of-the-art baselines. It maintains stable performance on long sequences (up to 1000 residues) and complex topologies, effectively addressing the length-dependent degradation issues observed in baseline diffusion models.
2) In binder design, to address the trade-off between design success rate and structural diversity, we introduce two inference modes: SeedProteo-R (Robust), which prioritizes longer secondary structure segments for higher design success, and SeedProteo-D (Diverse), which fosters topological richness. Experiments show robust performance across diverse targets, particularly excelling in challenging cases.
3) Furthermore, our analysis offers specific insights into how to achieve superior sequence-structure consistency within the atom14 representation schema [quPallatomUnlockingNew2025, butcher2025novo, starkUniversalBinderDesign]. We demonstrate that introducing a global MRF-based decoding module leads to more self-consistent outputs.
2 Method
2.1 Model Architecture
[…]
As shown in Figure 2, SeedProteo leverages the powerful non-equivariant atomic coordinate modeling capability of AF3-like folding model architectures, and therefore adopts a network structure consistent with such folding models: (i) an embedder module for feature initialization, (ii) an encoder composed of Pairformer blocks, and (iii) a diffusion module that denoises coordinates based on the encoder’s representations.
However, a key challenge arises because the folding model’s encoder is sequence-based. A trivial solution would be to replace the sequence with all [MASK] tokens, but such a degenerate biased input would render the encoder ineffective. To address this issue, SeedProteo additionally feeds noisy coordinates into the encoder to extract intrinsic geometric relationships: within the embedder, noisy coordinates are transformed into 1D sequence representation. This modification, however, introduces a computational cost—since the encoder, which contains cubic algorithmic complexity, becomes dependent on the noisy inputs. To mitigate the computational burden, we reduced the number of Pairformer layers from 48 to 12.
Figure 2: Overview of the SeedProteo framework. The right panel illustrates the input representation from a folding perspective. By modifying specific input channels within a nearly identical network architecture, we adapt the framework into the generative design model shown on the left.
2.2 Self-conditioning features
The distribution of protein atomic coordinates constitutes a specialized manifold. Although diffusion models can directly learn this coordinate distribution, a simple diffusion process lacking constraints often leads to the generation of structures that are geometrically plausible yet physically unrealizable. These generated backbones often exhibit poor designability, meaning that no amino acid sequence can stably fold into the generated conformation. Consequently, sequences derived from inverse folding on these structures typically fail to adopt the intended topology. Recent advances in protein design methodologies have validated the effectiveness of self-conditioning. We posit that employing effective self-conditioning features represents a promising pathway for substantially enhancing design capabilities. The following features are utilized as conditioning inputs in our model.
  1. 1.
    Sequence Sampling: SeedProteo employs a Markov Random Field (MRF)-based Sequence Module [renAccurateRobustProtein2024] to decode amino acid sequences. The probability of a sequence $\mathbf{x}$ is parameterized using the learned pair and single representations (Eq. 1):
    $\mathrm{P}(\mathbf{X}=\mathbf{x}\mid\mathbf{a},\mathbf{z})\propto\exp\bigg[\sum_{i=1}^{L}h_{i}(x_{i}\mid\mathbf{a}_{i})+\sum_{i=1}^{L}\sum_{j=i+1}^{L}e_{ij}(x_{i},x_{j}\mid\mathbf{z}_{ij})\bigg],$ (1)
    where $\mathbf{x}=(x_{1},\dots,x_{L})$ denotes the target amino acid sequence of length $L$, and $\mathbf{a},\mathbf{z}$ represent the learned single and pair representations, respectively. The term $h_{i}(\cdot)$ models the site-specific bias for amino acid $x_{i}$ conditioned on the single feature $\mathbf{a}_{i}$, while $e_{ij}(\cdot)$ captures the pairwise coupling propensity between residues $x_{i}$ and $x_{j}$ derived from the pair feature $\mathbf{z}_{ij}$ [balakrishnanLearningGenerativeModels2011]. We observe that this sequence-guided approach enhances the model’s co-design capability by mitigating potential conflicts between sequence and structural constraints. By integrating the MRF module, SeedProteo eliminates the need to infer amino acid types by translating atom14 coordinates. Instead, it directly utilizes the sequence inferred by the MRF, which incorporates semantic information to avoid misclassification of structurally similar (isosteric) amino acids, as the final output sequence.
  2. 2.
    Secondary Structure Sequence: SeedProteo accepts a secondary structure (SS) sequence [hekkelman2025dssp] composed of three canonical characters (H, E, L) and a special mask token (X). The model performs unmasking predictions based on the input masked SS sequence and subsequently utilizes the predicted SS sequence for self-conditioning (Eq. 2):
    $\mathcal{S}^{(0)}$ $=\mathcal{S}_{\text{input}}$ (2)
    $\mathcal{S}^{(t)}_{\text{pred}}$ $=f_{\theta}\left(\mathcal{S}^{(t-1)}_{\text{cond}}\right)$
    $\mathcal{S}^{(t)}_{\text{cond}}$ $=\alpha_{t}\cdot\mathcal{S}^{(t)}_{\text{pred}}+(1-\alpha_{t})\cdot\mathcal{S}^{(t-1)}_{\text{cond}}$,
    where $\mathcal{S}_{\text{input}}$ represents the initial, potentially masked secondary structure representation. At iteration $t$, $f_{\theta}(\cdot)$ denotes the neural network predictor parameterized by $\theta$, yielding the unmasked prediction $\mathcal{S}^{(t)}_{\text{pred}}$. This prediction is integrated into the conditioning state $\mathcal{S}^{(t)}_{\text{cond}}$ via a scalar gating factor $\alpha_{t}\in[0,1]$, which controls the update rate of the self-conditioning signal. This capability enables SeedProteo to process arbitrarily generated SS sequences, even those containing unknown segments, and adaptively refine them into a physically plausible SS sequence. We demonstrate that incorporating secondary structure features effectively guides the generation of de novo-like structures, significantly enhancing the success rate in binder design applications.
  3. 3.
    Structural Template: SeedProteo utilizes the denoised structure obtained from the previous step as a structural template for self-conditioning. Specifically, a binning operation is applied to convert the Cβ distance map into a one-hot tensor, which serves as a pairwise feature. We demonstrate that incorporating such template features helps stabilize the structural sampling trajectory.
2.3 Training Objective
SeedProteo employs a set of loss functions commonly used in AF3-like folding models [abramsonAccurateStructurePrediction2024], including a coordinate diffusion loss, a smooth LDDT loss, and a distogram loss. The coordinate diffusion loss is applied without pre-aligning the target and predicted structures, thereby compelling the model to learn equivariance directly through this loss. Additionally, a cross-entropy loss is used to supervise the MRF-sampled sequence and the predicted secondary structure sequence.
2.4 Training Pipeline
SeedProteo was trained from scratch. Inspired by the training strategies of folding models, we adopted a multi-stage training pipeline. In the first stage, the model was trained using a small crop size and strictly filtered monomer data. In the second stage, training continued with a larger crop size and an expanded dataset comprising comprehensively filtered monomer data. Meanwhile, motifs are incorporated during the training process with monomer data, where partial structures paired with corresponding sequences are provided to enhance the model’s responsiveness to conditional inputs. In the third stage, the large crop size was maintained, and training proceeded with a balanced 1:1 sampling ratio of monomer data and protein multimer data. Detailed descriptions of the datasets and hyperparameters are provided in (Appendix 5 and Table 5).
3 In-silico Results
We evaluated SeedProteo on two primary tasks: unconditional protein monomer generation and protein binder design. Unconditional generation refers to designing monomeric proteins starting from Gaussian noise, which serves to assess the model’s fundamental generative capability. Binder design, on the other hand, involves generating potential high-affinity binders based on the full atomic information of a given target protein, with the goal of enabling real-world applications.
Although numerous diffusion-based co-design methods are capable of jointly generating sequences and structures (in either backbone or all-atom form), existing results consistently indicate that the sequence consistency of their outputs is generally lower than that achieved by redesigning structures using ProteinMPNN [dauparasRobustDeepLearning2022]. Therefore, in the subsequent evaluations, for the structures generated by diffusion-based methods, we employed ProteinMPNN for sequence redesign to ensure fair and comparable assessments. Regarding the details of the evaluation procedure and the definitions of the metrics, please refer to Appendix 6.
Figure 3: Unconditional monomer benchmark. Stricter thresholds are applied to define designability (in Appendix 6).
3.1 Unconditional Generation Benchmark
3.1.1 Scalability and Structural Diversity
We first evaluate SeedProteo against state-of-the-art diffusion-based baselines, La-Proteina [geffnerLaProteinaAtomisticProtein2025], Proteina [geffnerProteinaScalingFlowbased2024], and RFDiffusion [watsonNovoDesignProtein2023], on the task of unconditional monomer generation. We focus on two critical dimensions: scalability to long sequences and structural diversity. As illustrated in Figure 3 (Left), SeedProteo consistently achieves superior design success rates across the entire length spectrum (100–1000 residues). While competing methods exhibit a sharp performance degradation as sequence length increases, dropping to near-zero success rates beyond 600 residues, SeedProteo maintains robust performance, achieving a success rate of $60\%$ even at length 1000. This suggests that our method effectively models the long-range dependencies and global constraints inherent in large protein assemblies, a capability that current baselines lack. In addition to quality, SeedProteo excels in generative diversity. Figure 3 (Right) plots the number of unique structural clusters found within the successful samples. Notably, as protein length increases, baselines suffer from severe mode collapse (e.g., RFDiffusion and Proteina generate few to no unique clusters at long lengths). In contrast, SeedProteo sustains a high volume of unique clusters, indicating that it captures a richer and more generalizable distribution of the protein structure space.
Table 1: Unconditional Design Benchmark Performance across Structural Topologies. We compare SeedProteo against baselines on three different fold types: $\alpha$-helix bundles (HHH), Hybrid (HEL), and $\beta$-sheets (EEE). Metrics reported in order: Design Success Rate (Succ., $\uparrow$), Unique Success Clusters (Uniq., $\uparrow$), and Novelty (Nov., $\downarrow$). The best performance is highlighted in bold. Dash (–) indicates model failure or insufficient successful samples.
Panel A: Primarily $\alpha$-Helix (HHH) Structures
Length SeedProteo (Ours) La-Proteina Proteina RFDiffusion
Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$
100 100% 17 0.89 93% 27 0.86 88% 22 0.83 87% 19 0.89
200 96% 24 0.85 96% 60 0.80 73% 21 0.83 73% 8 0.85
300 95% 24 0.83 85% 47 0.80 51% 13 0.88
400 98% 32 0.79 61% 40 0.76 39% 14 0.88
500 90% 36 0.81 67% 43 0.74 33% 12 0.85
600 86% 38 0.79 43% 9 0.88 27% 6 0.89
700 86% 49 0.82 38% 6 0.83 24% 15 0.89
800 65% 39 0.79 41% 10 0.86 4% 2 0.96
900 66% 45 0.75 41% 7 0.88
1000 63% 41 0.77 31% 5 0.88
Panel B: Hybrid Helix-E-Loop (HEL) Structures
Length SeedProteo (Ours) La-Proteina Proteina RFDiffusion
Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$
100 97% 26 0.82 87% 27 0.86 87% 13 0.82 71% 12 0.85
200 92% 32 0.79 90% 7 0.81 55% 6 0.82 38% 22 0.80
300 81% 21 0.77 72% 8 0.83 15% 4 0.85 11% 6 0.78
400 79% 26 0.77 62% 3 0.79 12% 2 0.91 3% 3 0.80
500 69% 20 0.75 60% 3 0.83 5% 2 0.87 1% 1 0.76
600 57% 23 0.77 26% 12 0.85
700 72% 28 0.77 9% 6 0.88 9% 1 0.93
800 56% 22 0.79 11% 6 0.86
900 44% 14 0.78 3% 2 0.89
1000 57% 20 0.74
Panel C: Primarily $\beta$-Sheet (EEE) Structures
Length SeedProteo (Ours) La-Proteina Proteina RFDiffusion
Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$ Succ. $\uparrow$ Uniq. $\uparrow$ Nov. $\downarrow$
100 100% 12 0.80 100% 1 0.88 50% 3 0.84
200 86% 6 0.80 100% 2 0.84 20% 1 0.75
300 80% 4 0.80 25% 1 0.85
400 73% 7 0.80 50% 1 0.85 57% 2 0.85
500 50% 1 0.68 50% 1 0.74
600 25% 4 0.86
700 8% 1 0.86
Note: Only lengths up to 700 are shown for EEE as all methods failed to generate valid long $\beta$-sheet structures.
Table 2: Structural Novelty ($\downarrow$) in Binder Design Benchmark. Transposed comparison of average novelty scores. We compare SeedProteo variants against baselines across ten targets. Lower is Better. Best performance per target is highlighted in bold, and the second best is underlined.
Method TrkA PD-L1 Insulin BHRF1 IL-7RA SC2RBD VEGF-A H1 IL-17A TNFA
Ours
SeedProteo-D 0.829 0.832 0.837 0.822 0.840 0.819 0.836 0.823 0.806 0.870
SeedProteo-R 0.905 0.913 0.911 0.872 0.917 0.858 0.901 0.890 0.855 -
Baselines
BindCraft 0.849 0.856 0.864 0.847 0.861 0.863 0.850 0.830 0.818
PXDesign 0.914 0.929 0.928 0.924 0.928 0.917 0.913 0.888 0.906
BoltzGen 0.908 0.924 0.929 0.928 0.885 0.915 0.902 0.885 0.863
RFDiffusion 0.932 0.934 0.927 0.946 0.916 0.912 0.940 0.938
RFDiffusion3 0.808 0.834 0.876 0.845 0.840 0.930 0.800
3.1.2 Fine-grained Evaluation across Structural Topologies
Previous generative models often exhibit a strong preference for simple $\alpha$-helical architectures [watsonNovoDesignProtein2023, quPallatomUnlockingNew2025], frequently failing to generalize to complex topologies such as all-$\beta$ sheets. Our empirical observations confirm that designability is highly topology-dependent: primarily $\beta$-sheet structures consistently yield lower success rates compared to helical bundles across all models. To dissect this phenomenon, we present a fine-grained evaluation in Table 1, breaking down performance by three distinct topologies: primarily $\alpha$-helix (HHH), primarily $\beta$-sheet (EEE) and hybrid Helix-Sheet-Loop (HEL). The results reveal a distinct advantage for SeedProteo in handling complex folds:
  • Complex Topology Robustness: In the Hybrid (HEL) category, SeedProteo achieves a 57% success rate at length 1000, whereas baselines largely fail to produce valid structures beyond length 600.
  • $\beta$-Sheet Generation: In the $\beta$-sheet (EEE) regime, which is notoriously difficult for generative models, SeedProteo is the only method capable of generating diverse, valid and novel structures (Success $50\%$ and Novelty <0.8<0.8) at lengths up to 500. Other models suffer from complete generation failure or mode collapse in this category.
  • Novelty and Diversity: Across all topologies, SeedProteo consistently achieves higher structural diversity (Unique Clusters) and favorable novelty scores, indicating that it does not merely retrieve training templates but effectively explores novel regions of the protein manifold.
3.2 Binder Design Benchmark
We benchmark SeedProteo against five leading open-source binder design methods: RFDiffusion [watsonNovoDesignProtein2023], RFDiffusion3 [butcher2025novo], BoltzGen [starkUniversalBinderDesign], PXDesign [ren2025pxdesign], and BindCraft [pacesaOneshotDesignFunctional2025]. Following the protocol established in AlphaProteo [zambaldi2024novo], we generate approximately 1,000 candidates for each of the 10 benchmark targets, uniformly covering the target-specific binder length ranges. For structure-generating diffusion models (SeedProteo, RFDiffusion variants, PXDesign, and BoltzGen), binder sequences are redesigned using ProteinMPNN. A notable exception is BindCraft, a hallucination-based approach [anishchenkoNovoProteinDesign2021], which optimizes sequences directly via gradient descent on folding metrics (e.g., ipTM). All designs are subsequently evaluated in silico using SeedFold [seedfold]. We adopt the success criteria defined in AlphaProteo: minimum inter-chain Predicted Aligned Error (PAE) $\leq 1.5$, binder pTM $\geq 0.8$, and complex RMSD <2.5<2.5Å.
In SeedProteo, we introduce two distinct Secondary Structure (SS) sampling modes to balance the trade-off between designability and structural diversity:
  • SeedProteo-R: This is a SS robust mode. Prioritizes fewer but longer continuous secondary structure segments. This mode imposes stronger structural stability constraints, typically yielding a higher design success number.
  • SeedProteo-D: This is a SS diverse mode. Allows for a higher frequency of shorter secondary structure segments. This mode encourages more complex topologies, resulting in greater structural diversity at the cost of slightly reduced raw success number.
Performance Analysis.
[…]
Beyond design success, structural diversity is a critical metric for downstream applicability. Figure 1 highlights that SeedProteo-D excels in generating structurally distinct binders, achieving the highest number of unique success clusters in 8 out of the 10 targets. Complementing this structural diversity, Table 2 demonstrates that SeedProteo-D also achieves the best novelty scores for successfully designed binders in 8 out of the 10 targets. Notably, it consistently outperforms all baselines, except RFDiffusion3, across every target; for example, while baselines like PXDesign often yield scores above 0.9, SeedProteo-D maintains superior novelty with scores in the 0.8 range. This indicates that SeedProteo does not merely converge on a single solution but effectively explores the structural landscape to provide a diverse portfolio of high-quality and novel candidates. Figure 4 presents case studies on challenging targets.
Figure 4: Binders generated by SeedProteo for the challenging multi-chain targets H1 (dimer), VEGF-A (dimer), and TNF-$\alpha$ (trimer). All displayed binders, colored in purple, meet the in silico success criteria.
Table 3: Benchmarking Co-design Capabilities of Diffusion-based Models for Binder Design.
[…]
Baselines
BoltzGen 627 0 10 17 21 99 1 1 5 0
RFDiffusion3 9 0 0 1 0 0 0 0 0 0
SeedProteo (Ours)
SeedProteo-R 296 92 100 380 232 303 16 127 47 1
SeedProteo-D 139 80 52 265 143 181 25 45 17 3
SeedProteo-M 133 77 25 154 143 189 12 47 13 1
Table 4: Ablation Study on Sequence Decoding Strategies.
Length Decoding Method scRMSD $\downarrow$ scTM $\uparrow$ pLDDT $\uparrow$ Succ. Rate $\uparrow$
100 Baseline (atom14) 1.11 0.94 88.56 0.90
SeedProteo (MRF) 1.63 0.93 88.65 0.84
200 Baseline (atom14) 3.87 0.86 78.06 0.68
SeedProteo (MRF) 2.08 0.92 84.36 0.80
300 Baseline (atom14) 4.71 0.83 73.40 0.52
SeedProteo (MRF) 2.30 0.91 80.33 0.76
3.3 Discussion on Co-Design Strategies
We here discuss the significance of co-design and methods for enhancing co-design capabilities. While two-stage pipelines (backbone generation followed by inverse folding) have dominated recent protein design benchmarks, our analysis suggests that simultaneous sequence-structure generation (co-design) offers potential advantages for complex targets.
Beyond the Inverse Folding Bottleneck
Empirical results (Figure 1 and Table 3) indicate that, on average, sequences derived from ProteinMPNN achieve higher success rates than those generated via pure co-design. This is expected, as ProteinMPNN is highly optimized for fixed-backbone recovery. However, we argue that reliance on inverse folding alone is not the ultimate solution for de novo binder design.
[…]
Decoding Mechanisms
We further compare the co-design mechanisms of SeedProteo with diffusion-based baselines, including RFDiffusion3 and BoltzGen, while excluding PXDesign and RFDiffusion as they are limited to backbone design rather than all-atom modeling. Both RFDiffusion3 and BoltzGen utilize variants of the atom14 schema, relying on removing specific virtual atoms to implicitly decode amino acid types. In contrast, SeedProteo employs atom14 strictly as a geometric input representation, while the final sequence identity is inferred via a Markov Random Field (MRF) module. This allows our model to capture higher-order residue dependencies that simple geometric decoding might miss.
3.4 Ablation Study
Impact of Secondary Structure Constraints
To evaluate the specific contribution of secondary structure (SS) constraints, we compare our approach against a baseline termed SeedProteo-M, which performs binder design using fully masked SS sequences as condition. As shown in Table 3, incorporating secondary structure (SS) sequences as conditioning constraints significantly enhances design success. This constraint effectively prunes the search space, guiding the diffusion process toward de novo-like fold topologies.
Impact of MRF module
To validate the effectiveness of our sequence decoding module, we conducted a supplementary ablation study on unconditional generation (Table 4). We established a baseline that predicts amino acid identities by aggregating features from the atom14 representation at each residue position, a strategy adopted by the recent method [quPallatomUnlockingNew2025].
As shown in Table 4, while the baseline achieves competitive performance on short sequences (Length 100), its capability degrades significantly as protein size increases. We reason this divergence to the inherent limitations of the decoding receptive fields. The atom14 schema relies on recognizing amino acid types via local side-chain geometry (implicit conformational recognition). However, the underlying AF3-based architecture primarily models pairwise inductive biases and may lack the capacity to capture long-range atomic dependencies solely through local atoms aggregation. Consequently, the baseline struggles to maintain global consistency in larger systems. Conversely, the MRF module explicitly leverages global pairwise features during inference. By modeling the sequence distribution conditional on the entire backbone topology, the MRF ensures that the generated sequence is self-consistent with long-range structural constraints, thereby preventing the structural degeneration observed in the baseline at longer lengths.
4 Conclusion
In this work, we presented SeedProteo, a diffusion-based all-atom protein design model. Our empirical benchmarks demonstrate that SeedProteo exhibits superior robustness across long sequences and complex topologies in unconditional design. Furthermore, it achieves state-of-the-art performance among open-source methods for binder design, particularly on hard and multi-chain targets. While effective co-design remains an open challenge in the field, we offer specific insights into achieving superior sequence-structure consistency via a global MRF module. We hope these findings will inspire significant community efforts to further advance the frontier of simultaneous sequence-structure generation.
Contributions
Project Lead
Wei Qu1
Contributors
Yiming Ma1,2,†, Fei Ye1, Chan Lu1, Yi Zhou1, Kexin Zhang1,3,†, Lan Wang1, Minrui Gui1,4,†
Work is done during their internship at Bytedance Seed.
Overall Technical Lead
Quanquan Gu1
Affiliation
1 ByteDance Seed
2 Peking University
3 ShanghaiTech University
4 University of California, Los Angeles
Acknowledgments
We thank Liang Hong, Zaixiang Zheng, Xinyou Wang, Jiasheng Ye, Jing Yuan, Yilai Li, Zhenghua Wang, Yuning Shen, Cheng-Yen Hsieh, Huizhuo Yuan, as well as other colleagues at ByteDance for their discussions and support.
[…]
5 Data Preparation and Curation
Training a foundation model for protein design, particularly one capable of binder generation, necessitates a diverse corpus comprising both monomeric and multimeric structures. While predicted monomer databases such as the AlphaFold Database (AFDB) [varadiAlphaFoldProteinStructure2022a] and ESMAtlas [lin2023evolutionary] provide extensive coverage of single-chain structures, existing multimer datasets derived from the RCSB PDB [berman2000protein] are limited in scale and often unsuitable for the specific two-chain requirements of binder design tasks. To address this, we constructed a composite training dataset by integrating curated monomer data with multimeric interactions sourced from Pinder [kovtun2024pinder] and HumanPPI [zhangPredictingProteinproteinInteractions2025]. We detail the processing and filtering pipelines for each category below.
5.1 Monomer Data Curation
Our monomeric dataset is derived exclusively from AFDB and ESMAtlas. For the ESMAtlas component, to mitigate potential quality issues associated with ESMFold [lin2023evolutionary] predictions, we retrieved the corresponding sequences from the MGnify database [richardson2023mgnify] and re-predicted their structures using AlphaFold2 (AF2) [jumper2021highly].
To ensure high data quality, we applied the following inclusion criteria to all monomer structures:
  • Sequence length between 50 and 768 residues.
  • Average pLDDT score >80>80.
  • Secondary structure coil fraction $50\%$.
To eliminate redundancy, we performed clustering using Foldseek [vankempenFastAccurateProtein2024] and MMseqs2 [steineggerMMseqs2EnablesSensitive2017] and retained only the cluster representatives (centroids). The final curated monomer dataset consists of approximately 0.5 million structures. Sequences that can be folded by SeedFold [seedfold] in a single-sequence setting are exclusively allocated for the training of the initial phase.
5.2 Multimer Data Curation
Protein-Protein Interactions (PPIs) are fundamental to training a robust binder design model. We sourced multimer data from two primary streams to cover experimentally determined interactions and high-quality predicted domain interactions.
Experimental PPIs (Pinder).
To align our training data with the standard inference scenario, where a binder chain is designed for a specific target chain, we utilized the Pinder dataset. Using the curated holo-structure PDB IDs and chain IDs provided by Pinder, we extracted relevant substructures from RCSB biological assemblies. We filtered these two-chain complexes by requiring: (1) a coil fraction of $50\%$ for each chain, and (2) a minimum interfacial Cβ distance of $8\text{\AA }$ to ensure meaningful contact. Following structural clustering via Foldseek-Multimer, we retained approximately 50,000 unique cluster representatives.
Augmented DDI Dataset (HumanPPI).
To overcome the scarcity of experimental PPI data, we augmented our training set with domain-domain interactions (DDIs). This approach relies on the established hypothesis that intrachain DDI interfaces resemble interchain PPI interfaces in terms of coevolutionary and physicochemical properties [lau2024exploring]. We utilized the HumanPPI database, which distills DDIs from accurately predicted monomer structures in the AFDB. We applied the identical filtering and clustering pipeline used for the Pinder dataset (coil check and distance constraints). This augmentation yielded an additional $\sim 0.1$ million interaction pairs.
Table 5: Training Stages.
Hyperparameter Stage 1 (Initial) Stage 2 (Fine-tuning 1) Stage 3 (Fine-tuning 2)
Crop Size 384 768 768
Batch Size 128 64 64
Motif Percentage 0% 20% 20%
Training Steps 50K 20K 30K
Data Sampling Distribution
    Strictly Filtered Monomer 100% 20% 10%
    Expanded Monomer Data 80% 40%
    Multimer Data 50%
6 Evaluation Settings
6.1 Unconditional Generation Evaluation
[…]
Topology-Aware Analysis.
To address the long-standing challenge in protein generative models, specifically the bias toward $\alpha$-helical structures and the difficulty in generating valid $\beta$-sheets, we incorporate a fine-grained analysis based on secondary structure topology. Based on secondary structure proportions, we classify the generated proteins into three distinct fold types. We define EEE (primarily $\beta$-sheet) as structures with over 40% sheet and under 20% helix content, and HHH (primarily $\alpha$-helical) as those with over 50% helix and under 10% sheet content; both categories require a loop fraction below 45%. All remaining structures are categorized as hybrid Helix-E-Loop (HEL).
6.2 Binder Design Evaluation
We evaluate binder design performance on a benchmark set of 10 targets, consistent with the validation set used in AlphaProteo (see Table 6). To rigorously assess the upper bound of design capability across different length scales, we adopt a dense sampling strategy: for each target, we sample 100 candidates at every 5-residue interval within the target-specific binder length range. For example, for the SC2RBD target (length range 80–120 residues), this results in a total of 900 sampled candidates.
For sequence recovery, we generate two sequences for each designed binder backbone using SolubleMPNN at a sampling temperature of $\tau=0.001$. The designs are then validated using SeedFold in single sequence with target template mode. A binder design is considered successful if it meets the following criteria:
  • Interface Predicted Aligned Error (min PAE interaction) <1.5<1.5;
  • Binder pTM score >0.8>0.8;
  • Complex RMSD <2.5<2.5Å.
Table 6: Benchmark Dataset for Binder Design. We evaluated performance on 10 diverse targets. The table lists the target details, including the PDB ID, the specific chains/residues used as the target region, the key hotspot residues defining the binding site, and the range of binder lengths sampled during inference.
Target PDB ID Target Region Hotspot Residues Length Range
BHRF1 2wh6 Chain A: 2-158 A65, A74, A77, A82, A85, A93 80-120
SC2RBD 6m0j Chain E: 333-526 E485, E489, E494, E500, E505 80-120
IL-7RA 3di3 Chain B: 17-209 B58, B80, B139 50-120
PD-L1 5o45 Chain A: 17-132 A56, A115, A123 50-120
TrkA 1www Chain X: 282-382 X294, X296, X333 50-120
Insulin 4zxb Chain E: 6-155 E64, E88, E96 40-120
H1 5vli Chain A: 1-50, 76-80, 107-111, 258-322; Chain B: 1-68, 80-170 B21, B45, B52 40-120
VEGF-A 1bj1 Chain V: 14-107; Chain W: 14-107 W81, W83, W91 50-140
IL-17A 4hsa Chain A: 17-131; Chain B: 19-127 A94, A116, B67 50-140
[…]
7 Baseline Sampling
In this section, we detail the specific configurations, checkpoints, and codebases used for all baseline methods included in our benchmarks.
7.1 Unconditional Benchmark
La-Proteina. We utilized the official implementation and checkpoints provided in the La-Proteina repository [^1] https://github.com/NVIDIA-Digital-Bio/la-proteina. To ensure optimal performance across different scales, we selected checkpoints based on the target sequence length: the LD1_ucond_notri_512.ckpt checkpoint was used for lengths 100–500, while LD3_ucond_notri_800.ckpt was employed for lengths 600–1000. All sampling was conducted using the default inference configurations.
Proteina. We used the code and checkpoints from the official Proteina repository [^2] https://github.com/NVIDIA-Digital-Bio/proteina. Similar to La-Proteina, we adopted a length-dependent checkpoint strategy: proteina_v1.1_DFS_200M_tir.ckpt was used for lengths 100–500, and proteina_v1.6_DFS_200M_notri_long_chain_generation.ckpt was used for lengths 600–1000. Default sampling parameters were applied in all cases.
RFDiffusion. We employed the official codebase available at the RFDiffusion repository [^3] https://github.com/RosettaCommons/RFdiffusion. All unconditional generation tasks were performed using the standard default sampling configurations provided by the authors.
7.2 Binder Design Benchmark
For the binder design task, all methods utilize identical target structures (in mmCIF format) and are conditioned on the exact same set of hotspot residues. For each target, we generate an equal number of candidate proteins with consistent lengths across all models. To ensure a fair comparison of generative capabilities, we disable any built-in auxiliary modules—such as pre-filtering, ranking, or early-stopping strategies—thereby restricting the evaluation strictly to the core sampling function.
BindCraft. We utilized the official implementation from the BindCraft repository [^4] https://github.com/martinpacesa/BindCraft. For binder generation, we adopted the default_4stage_multimer configuration under the advanced setting. To ensure a fair evaluation of the generative capability without post-hoc selection bias, we disabled both the filtering stage and the early stopping strategy. This ensured that every sampling attempt was completed and all raw outputs were retained, regardless of their internal quality metrics.
BoltzGen. We performed evaluations using the code provided in the BoltzGen repository [^5] https://github.com/HannesStark/boltzgen, specifically utilizing commit SHA 58c1eed. We employed the protein-anything protocol for inference. To bypass internal ranking and filtering mechanisms, we set the inference budget equal to the number of desired designs, ensuring a direct one-to-one output.
RFDiffusion. For binder design tasks, we set both noise_scale_ca and noise_scale_frame to 0 during sampling. The authors suggest that this deterministic setting yields superior performance for specific binder design scenarios.
RFDiffusion3. We utilized the implementation available in the RosettaCommons Foundry repository [^6] https://github.com/RosettaCommons/foundry, specifically commit SHA 7f6656e. We set align_trajectory_structures=False to prevent potential runtime errors during trajectory processing; all other configurations were maintained at default values.
PXDesign. We utilized the implementation in the PXDesign repository [^7] https://github.com/bytedance/PXDesign, specifically commit SHA ec6615c. We perform binder design for each target of a given hotspot using Generation Only mode.

### v2

@@
 Extensive benchmarks highlight the model’s capabilities across two distinct tasks: in unconditional generation, SeedProteo exhibits superior length generalization and structural diversity, maintaining robustness for long sequences and complex topologies; in binder design, it achieves state-of-the-art performance among open-source methods, attaining the highest in-silico design success rates, structural diversity and novelty.
+Finally, we validate SeedProteo through wet-lab assays on two therapeutic targets, achieving hit rates of 70%–80% and picomolar-level binding affinities, establishing leading results.
+To facilitate community adoption, we provide public access to SeedProteo via a webserver (https://seedfold.io/proteinDesign).
 […]
 […]
-Figure 1: Comparison of binder design success and diversity across ten target proteins. We benchmark exclusively against open-source methods using SeedFold in-silico metrics.
 1 Introduction
@@
 By integrating high-order couplings among distinct positions, we are able to sample globally feasible sequences and prevent being trapped in the local minima of simple patterns constituted by high-frequency amino acids.
-We conduct experiments on two tasks: unconditional sampling and binder design.
+We conduct extensive experiments on unconditional sampling, binder design, and in vitro validation.
 The results demonstrate that:
@@
 We demonstrate that introducing a global MRF-based decoding module leads to more self-consistent outputs.
+4) Beyond in-silico evaluation, we further validate SeedProteo with wet-lab experiments, demonstrating that designed binders exhibit strong binding in SPR assays with high hit rates across multiple therapeutic targets.
 2 Method
@@
   1. 1.
-    Sequence Sampling: SeedProteo employs a Markov Random Field (MRF)-based Sequence Module [renAccurateRobustProtein2024] to decode amino acid sequences.
+    Sequence Sampling: SeedProteo employs a Markov Random Field (MRF)-based Sequence Module [renAccurateRobustProtein2024, meynardfast] to decode amino acid sequences.
 The probability of a sequence $\mathbf{x}$ is parameterized using the learned pair and single representations (Eq.
@@
 Beyond design success, structural diversity is a critical metric for downstream applicability.
-Figure 1 highlights that SeedProteo-D excels in generating structurally distinct binders, achieving the highest number of unique success clusters in 8 out of the 10 targets.
+Figure SeedProteo: Accurate De Novo All-Atom Design of Protein Binders highlights that SeedProteo-D excels in generating structurally distinct binders, achieving the highest number of unique success clusters in 8 out of the 10 targets.
 Complementing this structural diversity, Table 2 demonstrates that SeedProteo-D also achieves the best novelty scores for successfully designed binders in 8 out of the 10 targets.
@@
 Beyond the Inverse Folding Bottleneck
-Empirical results (Figure 1 and Table 3) indicate that, on average, sequences derived from ProteinMPNN achieve higher success rates than those generated via pure co-design.
+Empirical results (Figure SeedProteo: Accurate De Novo All-Atom Design of Protein Binders and Table 3) indicate that, on average, sequences derived from ProteinMPNN achieve higher success rates than those generated via pure co-design.
 This is expected, as ProteinMPNN is highly optimized for fixed-backbone recovery.
@@
 By modeling the sequence distribution conditional on the entire backbone topology, the MRF ensures that the generated sequence is self-consistent with long-range structural constraints, thereby preventing the structural degeneration observed in the baseline at longer lengths.
+4 In Vitro Experiments
+Figure 4: Wet-lab validation of SeedProteo-designed binders.
+For each target (PD-L1, top; SC2RBD, bottom), three representative binding complex structures are shown alongside the SPR sensorgram of the highest-affinity binder.
+To validate the practical utility of SeedProteo beyond in-silico metrics, we conducted wet-lab experiments on two therapeutically relevant protein targets: the SARS-CoV-2 receptor-binding domain (SC2RBD) [cao2020novo] and Programmed Death-Ligand 1 (PD-L1) [yang2025design].
+These targets were selected due to their high clinical relevance—SC2RBD is a primary target for COVID-19 therapeutic antibody development, while PD-L1 is a critical immune checkpoint widely pursued in cancer immunotherapy—as well as their structural diversity, which allows us to assess the generalizability of our model across distinct binding interfaces.
+For each target, we generated several thousand minibinders using SeedProteo.
+The structures passing the in-silico criteria were further clustered using Foldseek to ensure structural diversity, from which 10 representative minibinders were selected per target for experimental characterization.
+The selected minibinders were recombinantly expressed in E.
+coli with an N-terminal His-SUMO tag and purified via Ni-NTA affinity chromatography followed by size-exclusion chromatography (detailed protocols in Appendix 10.1).
+Binding properties were characterized using surface plasmon resonance (SPR) on a Biacore 8K+ instrument, where candidates were first screened at two concentrations and then subjected to kinetic affinity measurements using either single-cycle or multi-cycle protocols (Appendix 10.2).
+Protein expression and SPR assays for all designed candidates were performed by WuXi AppTec (China).
+As shown in Figure 4, SeedProteo achieved hit rates of 70%–80% across both targets, with the best binders reaching picomolar affinity for PD-L1 and sub-nanomolar affinity for SC2RBD.
+These results demonstrate that our in-silico design pipeline reliably produces binders with strong experimental binding, bridging the gap between computational prediction and wet-lab validation.
 5 Conclusion
@@
 While effective co-design remains an open challenge in the field, we offer specific insights into achieving superior sequence-structure consistency via a global MRF module.
+Crucially, our wet-lab validation demonstrates that these in-silico advantages carry over to real-world efficacy, yielding experimentally confirmed binders with high hit rates and picomolar-level affinities.
 We hope these findings will inspire significant community efforts to further advance the frontier of simultaneous sequence-structure generation.
+6 Revision History
+Version 2, February 24, 2026
+  • Added in vitro experimental validation results.
+    • Added wet-lab experiments results with SPR binding characterization on SC2RBD and PD-L1 targets.
+    • Added revision history section.
+  • Updated several citations.
+Version 1, December 31, 2025
+Initial manuscript release.
 Contributions
@@
 We thank Liang Hong, Zaixiang Zheng, Xinyou Wang, Jiasheng Ye, Jing Yuan, Yilai Li, Zhenghua Wang, Yuning Shen, Cheng-Yen Hsieh, Huizhuo Yuan, as well as other colleagues at ByteDance for their discussions and support.
-[…]
-5 Data Preparation and Curation
-Training a foundation model for protein design, particularly one capable of binder generation, necessitates a diverse corpus comprising both monomeric and multimeric structures.
-While predicted monomer databases such as the AlphaFold Database (AFDB) [varadiAlphaFoldProteinStructure2022a] and ESMAtlas [lin2023evolutionary] provide extensive coverage of single-chain structures, existing multimer datasets derived from the RCSB PDB [berman2000protein] are limited in scale and often unsuitable for the specific two-chain requirements of binder design tasks.
-To address this, we constructed a composite training dataset by integrating curated monomer data with multimeric interactions sourced from Pinder [kovtun2024pinder] and HumanPPI [zhangPredictingProteinproteinInteractions2025].
-We detail the processing and filtering pipelines for each category below.
-5.1 Monomer Data Curation
-Our monomeric dataset is derived exclusively from AFDB and ESMAtlas.
-For the ESMAtlas component, to mitigate potential quality issues associated with ESMFold [lin2023evolutionary] predictions, we retrieved the corresponding sequences from the MGnify database [richardson2023mgnify] and re-predicted their structures using AlphaFold2 (AF2) [jumper2021highly].
-To ensure high data quality, we applied the following inclusion criteria to all monomer structures:
-  • Sequence length between 50 and 768 residues.
-  • Average pLDDT score >80>80.
-  • Secondary structure coil fraction $50\%$.
-To eliminate redundancy, we performed clustering using Foldseek [vankempenFastAccurateProtein2024] and MMseqs2 [steineggerMMseqs2EnablesSensitive2017] and retained only the cluster representatives (centroids).
-The final curated monomer dataset consists of approximately 0.5 million structures.
-Sequences that can be folded by SeedFold [seedfold] in a single-sequence setting are exclusively allocated for the training of the initial phase.
-5.2 Multimer Data Curation
-Protein-Protein Interactions (PPIs) are fundamental to training a robust binder design model.
-We sourced multimer data from two primary streams to cover experimentally determined interactions and high-quality predicted domain interactions.
-Experimental PPIs (Pinder).
-To align our training data with the standard inference scenario, where a binder chain is designed for a specific target chain, we utilized the Pinder dataset.
-Using the curated holo-structure PDB IDs and chain IDs provided by Pinder, we extracted relevant substructures from RCSB biological assemblies.
-We filtered these two-chain complexes by requiring: (1) a coil fraction of $50\%$ for each chain, and (2) a minimum interfacial Cβ distance of $8\text{\AA }$ to ensure meaningful contact.
-Following structural clustering via Foldseek-Multimer, we retained approximately 50,000 unique cluster representatives.
-Augmented DDI Dataset (HumanPPI).
-To overcome the scarcity of experimental PPI data, we augmented our training set with domain-domain interactions (DDIs).
-This approach relies on the established hypothesis that intrachain DDI interfaces resemble interchain PPI interfaces in terms of coevolutionary and physicochemical properties [lau2024exploring].
-We utilized the HumanPPI database, which distills DDIs from accurately predicted monomer structures in the AFDB.
-We applied the identical filtering and clustering pipeline used for the Pinder dataset (coil check and distance constraints).
-This augmentation yielded an additional $\sim 0.1$ million interaction pairs.
-Table 5: Training Stages.
-Hyperparameter Stage 1 (Initial) Stage 2 (Fine-tuning 1) Stage 3 (Fine-tuning 2)
-Crop Size 384 768 768
-Batch Size 128 64 64
-Motif Percentage 0% 20% 20%
-Training Steps 50K 20K 30K
-Data Sampling Distribution
-    Strictly Filtered Monomer 100% 20% 10%
-    Expanded Monomer Data 80% 40%
-    Multimer Data 50%
-6 Evaluation Settings
-6.1 Unconditional Generation Evaluation
-[…]
-Topology-Aware Analysis.
-To address the long-standing challenge in protein generative models, specifically the bias toward $\alpha$-helical structures and the difficulty in generating valid $\beta$-sheets, we incorporate a fine-grained analysis based on secondary structure topology.
-Based on secondary structure proportions, we classify the generated proteins into three distinct fold types.
-We define EEE (primarily $\beta$-sheet) as structures with over 40% sheet and under 20% helix content, and HHH (primarily $\alpha$-helical) as those with over 50% helix and under 10% sheet content; both categories require a loop fraction below 45%.
-All remaining structures are categorized as hybrid Helix-E-Loop (HEL).
-6.2 Binder Design Evaluation
-We evaluate binder design performance on a benchmark set of 10 targets, consistent with the validation set used in AlphaProteo (see Table 6).
-To rigorously assess the upper bound of design capability across different length scales, we adopt a dense sampling strategy: for each target, we sample 100 candidates at every 5-residue interval within the target-specific binder length range.
-For example, for the SC2RBD target (length range 80–120 residues), this results in a total of 900 sampled candidates.
-For sequence recovery, we generate two sequences for each designed binder backbone using SolubleMPNN at a sampling temperature of $\tau=0.001$.
-The designs are then validated using SeedFold in single sequence with target template mode.
-A binder design is considered successful if it meets the following criteria:
-  • Interface Predicted Aligned Error (min PAE interaction) <1.5<1.5;
-  • Binder pTM score >0.8>0.8;
-  • Complex RMSD <2.5<2.5Å.
-Table 6: Benchmark Dataset for Binder Design.
-We evaluated performance on 10 diverse targets.
-The table lists the target details, including the PDB ID, the specific chains/residues used as the target region, the key hotspot residues defining the binding site, and the range of binder lengths sampled during inference.
-Target PDB ID Target Region Hotspot Residues Length Range
-BHRF1 2wh6 Chain A: 2-158 A65, A74, A77, A82, A85, A93 80-120
-SC2RBD 6m0j Chain E: 333-526 E485, E489, E494, E500, E505 80-120
-IL-7RA 3di3 Chain B: 17-209 B58, B80, B139 50-120
-PD-L1 5o45 Chain A: 17-132 A56, A115, A123 50-120
-TrkA 1www Chain X: 282-382 X294, X296, X333 50-120
-Insulin 4zxb Chain E: 6-155 E64, E88, E96 40-120
-H1 5vli Chain A: 1-50, 76-80, 107-111, 258-322; Chain B: 1-68, 80-170 B21, B45, B52 40-120
-VEGF-A 1bj1 Chain V: 14-107; Chain W: 14-107 W81, W83, W91 50-140
-IL-17A 4hsa Chain A: 17-131; Chain B: 19-127 A94, A116, B67 50-140
-[…]
-7 Baseline Sampling
-In this section, we detail the specific configurations, checkpoints, and codebases used for all baseline methods included in our benchmarks.
-7.1 Unconditional Benchmark
-La-Proteina.
-We utilized the official implementation and checkpoints provided in the La-Proteina repository [^1] https://github.com/NVIDIA-Digital-Bio/la-proteina.
-To ensure optimal performance across different scales, we selected checkpoints based on the target sequence length: the LD1_ucond_notri_512.ckpt checkpoint was used for lengths 100–500, while LD3_ucond_notri_800.ckpt was employed for lengths 600–1000.
-All sampling was conducted using the default inference configurations.
-Proteina.
-We used the code and checkpoints from the official Proteina repository [^2] https://github.com/NVIDIA-Digital-Bio/proteina.
-Similar to La-Proteina, we adopted a length-dependent checkpoint strategy: proteina_v1.1_DFS_200M_tir.ckpt was used for lengths 100–500, and proteina_v1.6_DFS_200M_notri_long_chain_generation.ckpt was used for lengths 600–1000.
-Default sampling parameters were applied in all cases.
-RFDiffusion.
-We employed the official codebase available at the RFDiffusion repository [^3] https://github.com/RosettaCommons/RFdiffusion.
-All unconditional generation tasks were performed using the standard default sampling configurations provided by the authors.
-7.2 Binder Design Benchmark
-For the binder design task, all methods utilize identical target structures (in mmCIF format) and are conditioned on the exact same set of hotspot residues.
-For each target, we generate an equal number of candidate proteins with consistent lengths across all models.
-To ensure a fair comparison of generative capabilities, we disable any built-in auxiliary modules—such as pre-filtering, ranking, or early-stopping strategies—thereby restricting the evaluation strictly to the core sampling function.
-BindCraft.
-We utilized the official implementation from the BindCraft repository [^4] https://github.com/martinpacesa/BindCraft.
-For binder generation, we adopted the default_4stage_multimer configuration under the advanced setting.
-To ensure a fair evaluation of the generative capability without post-hoc selection bias, we disabled both the filtering stage and the early stopping strategy.
-This ensured that every sampling attempt was completed and all raw outputs were retained, regardless of their internal quality metrics.
-BoltzGen.
-We performed evaluations using the code provided in the BoltzGen repository [^5] https://github.com/HannesStark/boltzgen, specifically utilizing commit SHA 58c1eed.
-We employed the protein-anything protocol for inference.
-To bypass internal ranking and filtering mechanisms, we set the inference budget equal to the number of desired designs, ensuring a direct one-to-one output.
-RFDiffusion.
-For binder design tasks, we set both noise_scale_ca and noise_scale_frame to 0 during sampling.
-The authors suggest that this deterministic setting yields superior performance for specific binder design scenarios.
-RFDiffusion3.
-We utilized the implementation available in the RosettaCommons Foundry repository [^6] https://github.com/RosettaCommons/foundry, specifically commit SHA 7f6656e.
-We set align_trajectory_structures=False to prevent potential runtime errors during trajectory processing; all other configurations were maintained at default values.
-PXDesign.
-We utilized the implementation in the PXDesign repository [^7] https://github.com/bytedance/PXDesign, specifically commit SHA ec6615c.
-We perform binder design for each target of a given hotspot using Generation Only mode.